PLEASE HELP ME SOLVE THIS EQUATION WORD PROBLEM:

At the beginning of the school year , Jamie's height was 62 and 1/2 inches . During the year she grew 1 and 3/4 inches , 1 /8 of an inch more than she grew the previous year.
(A) What was jamie's height at the end of the school year?
(B) How tall was Jamie at the start of the previous school year?
Thank you

Answer by elima(1433)

a)You need to add her hieght at the beginning of the year to how much she grew.
62 1/2 + 1 3/4 = 62 2/4 + 1 3/4=62 5/4= 63 1/4
b)Now subtract 1/8 from 1 3/4,
1 3/4 - 1/8 = 1 5/8
Now subtract that from 62 1/2
62 1/2 - 1 5/8= 60 7/8
So her hieght at the beginningof last year was - 60 7/8
Hope you understand
=)